OpenStudy (anonymous):

@Cuanchi with calculate the amount of energy in joules required to change 10 grams of solid mercury at its melting point to mercury vapor at the boiling pt. the melting point, boiling point and specific heat capacity are -39C 375C and 0.140 J/gC, respectively. heat of fusion= 11.4 J/g heat of vaporization; 295 J/g. Compare with amount of heat needed to change 10 grams of ice at 0C to steam @100C

OpenStudy (anonymous):

@Cuanchi ppplz

OpenStudy (anonymous):

@Abhisar cant do this problem

OpenStudy (anonymous):

@ganeshie8 help pplz

OpenStudy (anonymous):

@lesliejcastorena help me

OpenStudy (anonymous):

@iamabatman help me

OpenStudy (anonymous):

@adamaero help me

OpenStudy (anonymous):

@sleepyjess i need help

OpenStudy (anonymous):

@sleepyjess u know how to do it

OpenStudy (anonymous):

@pplz idek how to do it

OpenStudy (anonymous):

@jfraser i need help

OpenStudy (anonymous):

@abmon98 i need help

OpenStudy (anonymous):

@chvmijay i need help NOW

OpenStudy (sleepyjess):

@ganeshie8

OpenStudy (cuanchi):

E=m Ce delta T you have to calculate the energy of the 10 g of Hg(s) to became liquid E1= 10g x11.4J/g= then the energy to heat it up the 10 g of Hg from -39C to 295C E2= m Ce delta T =10g x 0.140J/g x (295+39)= then the energy to convert the Hg(l) to gas E3= 10g x 295 J/g= finally you add E1+E2+E3 do the same for the water look in tables the heat of fusion and vaporization and Ce for H2O